Solve the system using matrices. If there is no solution or if there are infinitely many solutions and system's equations are dependent, so state.
4x-4y=12
7x-7y=21

Answer by MathLover1(20849)


4x-4y=12+....eq.1..both sides divide by 4
7x-7y=21...both sides divide by 7
---------------------------------------------------
x-y=3+...eq.1
x-y=3..eq.2
----------------------
eq.1 is same as ..eq.2, so, the two lines are identical (one lies perfectly on top of the other) and intersect at all points of both lines and there are an infinite number of solutions and the system is dependent